The sum of an integer and its square is 42. Find the integer.
Answer by checkley71(8403)

X+X^2=42
X^2+X-42=0
(X+7)(X-6)=0
X+7=0
X=-7 ANSWER.
X-6=0
X=6 ANSWER.